OpenStudy (anonymous):

how many atoms and molecules are present in 124 gram of phosperus

OpenStudy (abb0t):

First you need to divide by the molar mass of \(\sf \color{red}{phosphorus}\)*. Remember that molar mass has the units: \(\sf \color{blue}{\frac{grams}{mol}}\) You can find the molar mass of phosphorus, P, by looking at your periodic table. If you don't have one, you can see use this one: http://ptable.com/ So, here is the route to be taken: \(\sf given~mass (\cancel{g}) \times \frac{mol}{\cancel{grams}~of~1~P~atom}\) = mol of Phosphorus. Next, use avogadro's constant to solve for both atoms/molecules. This constant is: \(\sf \color{green}{6.022 \times 10^{23} \frac{atoms}{mol}}\) So, you have: \(\sf \cancel{mol}~of~P \times \frac{6.022 \times 10^{23}~atoms}{\cancel{mol}}\) = atoms or molecules. See how the units cancel out nicely for each step. You're final answer should be some very large number times 10\(^{24}\). Which makes sense. There should be a lot of molecules in a relatively large mass such as the one given.